# Take Your Best Shot: Sampling-Based Next-Best-View Planning for Autonomous Photography & Inspection

### v1

@@ IV-B Candidate View Evaluation
The criteria for an optimal inspection picture from the previous sections can be used to assess the quality of any viewpoint given the vehicle’s current knowledge of the environment. As assumed in Sec. III, the vehicle is equipped with a depth sensor (e.g. lidar, stereo camera, RGB-D) and is capable of updating the occupancy map of the environment $\mathcal{M}$. In this application, the occupancy map is updated using recursive Bayesian updates [16], although any method can be used. The occupancy map ensures that as the robot moves through the environment, obstacles and obstructions are stored and planned for in subsequent iterations using $\mathcal{M}$. In turn, this requires that a view be reevaluated, since the quality of a view may increase or decrease as new occupied areas are observed. Thus, the quality of a viewpoint must be assessed at runtime; however, determining the vehicle’s view of the target in a cluttered, partially-known environment is complex and potentially creates a discontinuous or highly non-linear solution space for (1). To reduce the complexity of the computation, we use ray-casting to determine the quality of a viewpoint.
Ray-casting is a classical technique in computer vision that allows complex interactions and quality estimations to be discretely modeled [17]. Similar to [18], we use ray-casting and (2) to estimate the quality of a vehicle’s view. In our approach, we use a geometric ray-sphere transformation to estimate interactions with obstacles and the target efficiently. The set of coordinates that terminate at the robot’s maximum observation range $d_{\text{max}}$ are derived from a set of unit vectors $\mathbf{U}$ that emanate from a viewpoint and are defined as follows:
@@
We also define the radius $c_{r}$ to check each ray for an intersection with the discretized target space $\mathbf{T}$ or an obstacle coordinate. As illustrated in Fig. 5, each coordinate of interest $\bm{z}$ (e.g., target or obstacle coordinate) is checked for intersection with each ray defined by the viewpoint and the set $\mathbf{U}$. We use a geometric approach to project the line segment between the viewpoint $\bm{p}$ and the coordinate of interest along the vector $\bm{\hat{u}}$ corresponding to the coordinate $\bm{x}_{s}$. Algorithm 1 gives an overview of the ray-tracing procedure for an example coordinate $\bm{z}$.
@@
Given that we can use this algorithm to find intersections with target coordinates and obstacles, we update each $\bm{x}_{s}\in\mathbf{X}_{s}$ using Algorithm 1 and store whether each $\bm{x}_{s}$ is associated with the target, logically denoted by $y_{s}\in Y_{s}$.
@@
Subsequently, we can determine the values of the variables in (3) and (4) that intersect the target. $\phi_{i\min}$ and $\phi_{i\max}$ are the minimum and maximum angles between the normal vector of the frame and the vectors that intersect the target along the $i^{\text{th}}$ dimension. In the same way, $\ell_{i1}$ and $\ell_{i2}$ are determined by the distance to the point of intersection between the target and the rays along the $i^{\text{th}}$ dimension. Using this methodology, we can similarly compute the score in (2), but in a discretized manner, increasing the computation speed for each candidate view in the environment.
@@ IV-C Target Coverage Estimation using GP Interpolation
where $f(\mathbf{X})$ is functionally interpreted in this application as whether a target coordinate $\bm{x}\in\mathbf{X}$ has been captured by the inspecting vehicle (i.e., a zero if the coordinate has not been observed and a one if the coordinate has been inspected before). For interpolation of the target space, the widely used radial basis function (RBF) kernel is employed to capture the spatial relationship between two points and likewise interpolate their value using the Gaussian process regression with the kernel equation formulated as:
@@
where $\sigma_{l}\in[0,1]$ and $\sigma_{f}\in[0,1]$ are two hyperparameters. In this work, we tuned the hyperparameters by minimizing the negative log-likelihood of the training data. In this application, the hyperparameters were tuned using PSO because of its flexibility, given that it could be adapted to a higher speed or higher accuracy depending on the convergence criteria [22]. Additionally, the variable $\mathbf{X}\in\mathbb{R}^{m\times n_{d}}$ is a set of $m$ sampled target coordinates from the ray-casting discussed in the previous section. We can estimate the target coverage as the difference between the prior target coverage $\mu(\mathbf{X})$ and the estimated target coverage using the posterior mean, $\mu^{*}(\mathbf{X})$, from a new viewpoint $\bm{p}$ given the set of sample target points $\mathbf{X}_{s}\in\mathbb{R}^{s\times n_{d}}$. As such, the computational complexity of the target coverage is significantly reduced by the new equation
@@
The variable $\sigma_{n}$ is a constant in this case representing sensor noise and $Y_{s}$ is the binary 1D matrix signifying if the sample terminal ray-cast coordinates in $\mathbf{X}_{s}$ were in range of a target coordinate. With this belief update, we estimate the probability of additional coverage of the target from a candidate viewpoint $\bm{p}$ and (2) can be efficiently computed for each candidate viewpoint $\bm{p}$. However, computing every candidate view in the environment is intractable at runtime, especially given that (2) may have a non-linear or discontinuous solution space. To overcome this limitation, we incorporate particle swarm optimization to dynamically determine the optimal viewpoint in real-time.
@@ IV-D Candidate View Particle Swarm Optimization
As previously stated in Sec. IV-C, PSO is useful for solving a highly non-linear problem and is adaptable to runtime considerations. In our inspection application, a candidate viewpoint is represented as a point in a $N$-dimensional solution space. A swarm of particles is used to represent potential viewpoints. Each particle, $\bm{p}_{i}(t)$, is the coordinate of the $i^{\text{th}}$ candidate view at time $t$, and its velocity components are represented as $\bm{v}_{i}(t)$. The coverage $G$ is calculated for each viewpoint and the particles migrate toward the particle with the highest evaluation score.
@@
where $\bm{b}_{i}(t)$ and $\bm{g}(t)$ are the global best position evaluated by the particles and the global best position of all particles, respectively. The acceleration coefficients $c_{1}$ and $c_{2}$ are non-negative constants, which weigh the influence of the personal and global bests in the search process. The inertia weight $w$ balances the local and global exploration of the particles in the search space [23]. To avoid divergence, the value of $w$ must be between 0 and less than 1. The coefficients $c_{1}$ and $c_{2}$ can be adjusted to influence the particles to explore the area around their local optima (when $c_{1}>c_{2}$) or the area around the global optima (when $c_{2}>c_{1}$). This can be changed during the heuristic process to promote exploration or exploitation. Variables $r_{1}$ and $r_{2}$ are random numbers between 0 and 1 and are used to perturb the particles to encourage exploration.
Fig. 7 shows the result of our approach in 3D in which the vehicle detects an obstacle that blocks the target and propagates particles given the updated occupancy map. The vehicle moves toward the best view, as defined by the best global position of all particles. As the vehicle moves, the particles iteratively migrate in search of the best viewpoint. Once the vehicle reaches the best viewpoint location, it will capture a photo and update the observed portions of the target as shown in Fig. 7. This allows the vehicle to move to unobserved portions of the target in future iterations. In this instance, the best view of the target was shifted due to the obstacle. Since the raycasting is uniform from the vehicles’ position, the right side of the target was estimated as unseen rather than fully reflecting the ground truth outcome. We suggest that this is a beneficial outcome, as the vehicle was able to travel to the estimated “unseen” part of the target to obtain an undistorted image of the other side rather than the small left-hand portion.
Figure 6: An example using GP interpolation to estimate the visibility of a target in a two-dimensional space. The color bar to the right of (a) displays the ground truth target coverage, while (b) illustrates the estimation of the coverage using GP interpolation.
Figure 7: An example of using GP to estimate the visibility of the target in 3-dimensional space. (a) presents the settings, (b) shows the estimated GP result from the ray hits. (c) provides the ground truth for visibility.
IV-E Comparison Discussion
We note that this is a novel approach for photographing and inspecting targets, minimizing the number of photos to capture and evaluating the quality of the images. Therefore, other methods can incorporate these scoring metrics as part of their planning approach. We demonstrate an example of such modification, where we compare a sampling-based frontier approach from [24] applied to a target space in Fig. 8. Fig. 8 shows that the typical sampling-based frontier approach will value efficient exploration and consider target areas explored. The frontier approach captures a new image each time a previously unseen portion of the target is uncovered. It takes a total of 25 photos, each of which fails to match the quality metrics defined in our work. In contrast, the proposed approach shown in Fig. 8 only takes 2 inspection photos that maximize the designed evaluation metrics.
Figure 8: Comparison of sampling-based frontier approach (a) with the proposed NBV-based approach (b).
@@ V Simulations
To demonstrate the robustness of the proposed method, we conducted extensive simulations using different vehicles equipped with sensors of varying capabilities in diverse environmental contexts. In these simulations, the vehicle operates in a partially-known environment defined by operational boundaries, target position, and dimensions. As stated in Sec. III, we assume that the vehicle is equipped with two primary sensor types: 1) a depth sensor for obstacle detection (e.g., RGB-D camera, Lidar, etc.), and 2) an optical camera used for photography. We ran all tests in MATLAB on a Lenovo ThinkPad P14s laptop with a 2.80 GHz Intel 4-Core i7 processor and 32 GB RAM.
In the 2D simulation, the vehicle operates in a 100m×$\times$100m grid world with a resolution of 1m and uses a hybrid A* for path planning, benefiting from its ability to rapidly compute optimal paths in 2D environments while accounting for vehicle dynamics. The objective is to inspect the four sides of a building measuring 30m×$\times$15m. Various shapes of unknown obstacles are placed in front of three sides of the building. Both the depth sensor and the camera have a field of view (FOV) of $\Phi=[-\frac{\pi}{4},\frac{\pi}{4}]$. The range of the RGB-D sensor is $r_{max}=25$m. We set $\beta=0.8$ and the utility functions are designed in the style of logistic functions:
@@
We choose logistic functions because their inherent property of rapid decline in value allows for the effective differentiation between preferred values and those less desirable. For the utility function $U_{d}$, a more substantial penalty should be applied as $x$ approaches $1$, indicating increased perspective distortion. Therefore, we select $\mathbf{q}_{d}=[0.3,0.7,20,-0.75]$. Regarding $U_{s}$, we construct a piece-wise function with $\mathbf{q}_{s}=[0,1,-20,-0.5]$ for $x\in[0,\beta]$ and $\mathbf{q}_{s}=[0,1,30,-1]$ for $x\in(\beta,1]$ to ensure the utility’s peak aligns with the desired proportion $\beta$. $U_{d}$ diminishes as $x$ diverges from $\beta$, encouraging $\beta$ proportion of the target within the frame while preventing the view from intersecting with the frame edges [1]. Users can also alter these functions or design functions in similar styles to fit their applications. The utility functions are not dependent on the target or environment and were used in all simulations and experiments.
We use PSO to search for the optimal viewpoint by uniformly distributing 20 particles across the environment and iteratively refining the search. Fig. 9 shows the top-down view of some keyframes during the simulation. The heatmap on the second row is generated by computing the expected target coverage and image quality for every viewpoint in the environment using (2). As illustrated in Fig. 9, the vehicle takes pictures from the best viewpoint found by the PSO on each side of the building and takes several pictures from different points of view for the same target when necessary. The L2 norm of the error between the global optimal and the results obtained through PSO is $0.8538\pm 0.6816m$. Considering the dimension of the target, this error is neglectable. It’s noteworthy that this error could be mitigated by either adding more particles or increasing the number of optimization iterations. GP estimates of the inspected portions are shown as colored bars on the target. The evaluation performance for this 2D example occurs at a rate of 1.79ms per candidate viewpoint.
A 3D simulation result is shown in Fig. 10 in which a $\times 1$m×$\times 10$m target is positioned in a $\times 30$m×$\times 40$m space with an obstacle placed in front of it. In this case, we employ $*{}^{*}$ for its speed and efficiency in computing a path in a complex 3D environment. The RGB-D sensor has identical horizontal and vertical FOV. As indicated in Fig. 10, the aerial vehicle routes to the optimal viewpoint from where it can see the majority of the target and avoids the view from being occluded by the obstacle. The corresponding heatmaps orthogonal at the viewpoint are provided to show the optimality of the viewpoint. The evaluation for the 3D example occurs at a rate of 5.48ms per candidate viewpoint.
Figure 9: A top-down perspective of a 2D simulation featuring a quadrotor assigned to inspect four sides of a building. The frames are chronologically arranged from left to right. The first row captures 1) the quadrotor states, 2) exploration progress, and 3) the inspected portion of the target. The heatmaps in the second row show the ground truth information gain at all viewpoints in the environment.
@@ VI Experiments
First, to test the effectiveness of our approach, the Tello was started from the corner of the room and tasked to inspect an object in a free space. From the left to the right in Fig. 11, we demonstrate: 1) the bird view of Tello taking a picture, 2) the GP result shows the entire target is inspected, 3) the heatmap shows the ground truth information gain each viewpoint in the environment (with the UAV at the optimal position), and 4) the target lays in the center and takes up the desired $80\%$ of the horizontal FOV in the final picture.
Fig. 12 shows the result for a case with obstacles blocking the direct view of the facade of a target object. Thanks to our approach, the UAV realizes that it needs to take two pictures from each side of the T-shape obstacle to cover the entire front surface of the object and it only needs to take one perspective shot to look over the obstacle and inspect the entire side surface.
@@
Fig. 13 shows results for a setting with a slash-shaped obstacle. This time the ROSbot UGV locates the best vantage point at which the obstacle appears as a thin line.
@@
To further reinforce these results, we performed virtual experiments using a simulated RotorS Firefly equipped with a stereo camera. This experiment ran on the same hardware in Sec. V in C++. The evaluation of each candidate viewpoint occurs in 2.39ms. Fig. 14 shows the results of an example virtual experiment.
Figure 14: An example virtual experiment where a tree obstructs the view of the target. (a) and (b) show the initial and final view. (c) and (d) show the trajectory followed to the best viewpoint in gazebo and rviz.
As shown in Fig. 14, the vehicle detects the tree in front of the house using the stereo camera and stores the readings in an OctoMap [25]. The vehicle uses particle swarm optimization to maximize the equation in (2) where the result is a viewpoint that maximizes the information gain while minimizing distortion. The picture in Fig. 14 shows the resulting image from this viewpoint. For this experiment, we set $\beta$ to 0.8 and use $*{}^{*}$ to route the vehicle around obstacles. The Gazebo snapshot in Fig. 14 shows the planned trajectory of the UAV. The Rviz snapshot in Fig. 14 shows the resulting plan given the observed obstacles, where the red-shaded region represents the target.

### v2

@@ IV-B Candidate View Evaluation
The quality of pictures taken at viewpoints is evaluated based on the vehicle’s knowledge of the environment. The vehicle updates the environment’s occupancy map $\mathcal{M}$ using recursive Bayesian updates [16], which helps re-evaluate viewpoints as it moves, as view quality can change with new observations. To manage the complexity of determining optimal views in cluttered environments, ray-casting is used to assess viewpoint quality and reduce computational load.
Ray-casting, a classical technique in computer vision, allows for discrete modeling of complex interactions and quality estimations [17]. Following [18], we use ray-casting and (1) to estimate the view quality. In our approach, a geometric ray-sphere transformation is employed to efficiently estimate interactions with obstacles and the target. The set of coordinates terminating at the vehicle’s maximum observation range $d_{\text{max}}$ is derived from a set of unit vectors $\mathbf{U}$, defined as:
@@
We also define a radius $c_{r}$ to check each ray for intersections with the discretized target space $\mathbf{T}$ or obstacle coordinates. As shown in Fig. 5, we assess the intersection of a coordinate of interest $\bm{z}$ (e.g., target or obstacle) with each ray from the viewpoint in the set $\mathbf{U}$. Using a geometric approach, we project the line segment from the viewpoint $\bm{p}$ to the coordinate along the vector $\bm{\hat{u}}$ for the coordinate $\bm{x}_{s}$. Algorithm 1 overviews the ray-tracing procedure for a coordinate $\bm{z}$.
@@
Given that this algorithm can find intersections with target coordinates and obstacles, each $\bm{x}_{s}\in\mathbf{X}_{s}$ is updated using Algorithm 1, and whether each $\bm{x}_{s}$ is associated with the target is stored, denoted logically by $y_{s}\in Y_{s}$.
@@
Subsequently, we can find the values in (2) and (3) that intersect the target. $\phi_{i\min}$ and $\phi_{i\max}$ represent the minimum and maximum angles between the normal vector and the vectors that intersect the target along the $i^{\text{th}}$ dimension. Similarly, $\ell_{i1}$ and $\ell_{i2}$ are determined by the distance to the intersection point along dimension $i^{\text{th}}$. This approach allows us to compute the score in (1) discretely, increasing the computation speed for each candidate view.
@@ IV-C Target Coverage Estimation using GP Interpolation
where $f(\mathbf{X})$ is functionally interpreted in this application as whether a target coordinate $\bm{x}\in\mathbf{X}$ has been captured by the robot (i.e., a zero if the coordinate has not been observed and a one if the coordinate has been captured before). For interpolation of the target space, the widely used radial basis function (RBF) kernel is employed to capture the spatial relationship between two points and likewise interpolate their value using the Gaussian process regression with the kernel equation formulated as:
@@
where $\sigma_{l}\in[0,1]$ and $\sigma_{f}\in[0,1]$ are two hyperparameters. In this work, we tuned the hyperparameters by minimizing the negative log-likelihood of the training data. In this application, the variable $\mathbf{X}\in\mathbb{R}^{m\times n_{d}}$ is a set of $m$ sampled target coordinates from the ray-casting discussed in the previous section. We can estimate the target coverage as the difference between the prior target coverage $\mu(\mathbf{X})$ and the estimated target coverage using the posterior mean, $\mu^{*}(\mathbf{X})$, from a new viewpoint $\bm{p}$ given the set of sample target points $\mathbf{X}_{s}\in\mathbb{R}^{s\times n_{d}}$. As such, the computational complexity of the target coverage is significantly reduced by the new equation
@@
The variable $\sigma_{n}$ is a constant in this case representing sensor noise and $Y_{s}$ is the binary 1D matrix signifying if the sample terminal ray-cast coordinates in $\mathbf{X}_{s}$ were in range of a target coordinate. With this belief update, we estimate the probability of additional coverage of the target from a candidate viewpoint $\bm{p}$ and (1) can be efficiently computed for each candidate viewpoint $\bm{p}$. Fig. 6 showcases the estimated view coverage using ray tracing and GP in 2D. However, computing every candidate view in the environment is intractable at runtime, especially given that (1) may have a non-linear or discontinuous solution space. To overcome this limitation, we incorporate particle swarm optimization to dynamically determine the optimal viewpoint in real-time.
@@ IV-D Candidate View Particle Swarm Optimization
In our autonomous photography framework, a candidate viewpoint is represented as a point in a $N$-dimensional solution space. A swarm of particles is used to represent potential viewpoints. Each particle, $\bm{p}_{i}(t)$, is the coordinate of the $i^{\text{th}}$ candidate view at time $t$, and its velocity components are represented as $\bm{v}_{i}(t)$. The coverage $G$ is calculated for each viewpoint and the particles migrate toward the particle with the highest evaluation score.
@@
where $\bm{b}_{i}(t)$ and $\bm{g}(t)$ are the global best position evaluated by the particles and the global best position of all particles, respectively. The acceleration coefficients $c_{1}$ and $c_{2}$ are non-negative constants, which weigh the influence of the personal and global bests in the search process. The inertia weight $w$ balances the local and global exploration of the particles in the search space [22]. To prevent divergence, $w$ must be between 0 and less than 1. Adjusting coefficients $c_{1}$ and $c_{2}$ influences particles to explore around local optima (when $c_{1}>c_{2}$) or global optima (when $c_{2}>c_{1}$). This can be modified during the heuristic process for exploration or exploitation. Variables $r_{1}$ and $r_{2}$, random numbers between 0 and 1, perturb particles to encourage exploration.
Fig. 7 illustrates our 3D approach where the vehicle detects an obstacle obstructing the target and propagates particles using the updated occupancy map. As the vehicle moves toward the best viewpoint, the particles iteratively search for the optimal view. Upon reaching the best viewpoint, the vehicle captures a photo and updates the observed target areas, as shown in Fig. 7. The obstacle shifts the optimal viewpoint, causing the right side to be marked as unseen and targeted for capture in future iterations.
Figure 6: An example of GP interpolation for 2D target visibility. The color bar (a) shows true target coverage and (b) shows GP interpolated coverage.
Figure 7: An example of using GP to estimate the visibility of the target in 3D space. (a) shows the environment, (b) shows the estimated GP result from the ray hits. (c) provides the ground truth visibility.
@@ V Simulations
To demonstrate the robustness of the proposed method, we conducted extensive simulations using different vehicles equipped with sensors of varying capabilities in diverse environmental contexts. In these MATLAB simulations, the vehicle operates in a partially-known environment defined by operational boundaries, target position, and dimensions. The vehicle is equipped with two primary sensor types: 1) a depth sensor for obstacle detection (e.g., RGB-D camera, Lidar, etc.), and 2) an optical camera.
In the 2D simulation, the vehicle operates in a 100m×$\times$100m grid world with a resolution of 1m and uses hybrid A* for path planning. The objective is to capture the four sides of a building measuring 30m×$\times$15m. Various shapes of unknown obstacles are placed in front of three sides of the building. Both the depth sensor and the camera have a field of view (FOV) of $\Phi=[-\frac{\pi}{4},\frac{\pi}{4}]$, and the range of the RGB-D sensor is $r_{max}=25$m.
We set $\beta=0.8$, and the utility functions are designed in the style of logistic functions:
@@
We chose logistic functions for their rapid decline, allowing effective differentiation between preferred values and less desirable ones. For the utility function $U_{d}$, which penalizes as $x$ approaches $1$, we select $\mathbf{q}_{d}=[0.3,0.7,20,-0.75]$. For $U_{s}$, we construct a piecewise function with $\mathbf{q}_{s}=[0,1,-20,-0.5]$ for $x\in[0,\beta]$ and $\mathbf{q}_{s}=[0,1,30,-1]$ for $x\in(\beta,1]$. These functions ensure the utility’s peak aligns with the desired proportion $\beta$, reducing distortion and preventing view intersection with the frame edges [1].
We employ Particle Swarm Optimization (PSO) with 20 particles to find the optimal viewpoint. Fig. 8 shows keyframes from the simulation and heatmaps display expected target coverage and image quality metrics from (1). The L2 norm error between the globally optimal solution and the PSO results is $0.8538\pm 0.6816$m, which is insignificant considering the size of the target.
Figure 8: A 2D simulation from a top-down view, The first row displays 1) quadrotor states, 2) exploration progress, and 3) inspected target portions. The second row shows the ground truth information gain for all viewpoints.
In the 3D simulation (Fig. 9), a target measuring 10m×$\times$1m×$\times$10m is positioned inside a 30m×$\times$30m×$\times$40m area that contains an obstacle in front. We adopt RRT for path planning due to its effectiveness in complex 3D spaces. The aerial vehicle navigates to the optimal viewpoint while avoiding obstacles. The 3D example evaluation is conducted at a rate of 5.48ms per candidate viewpoint.
@@
We note that this approach minimizes the number of photos needed and evaluates image quality using defined metrics. Fig. 10 illustrates a comparison between our method and a sampling-based frontier approach [23]. While the frontier method captures 26 photos to cover the entire target, none meet our quality criteria. In contrast, our approach captures only 2 photos that capture the whole target maximizing the evaluation metrics.
Figure 10: Comparison of the proposed approach (a) with the sampling-based frontier approach (b).
@@ VI Experiments
First, to test the effectiveness of our approach, the Tello was started from the corner of the room and tasked to inspect an object in a free space. From the left to the right in Fig. 11, we show: 1) the bird view of Tello taking a picture, 2) the GP result shows the entire target is inspected, 3) the heatmap shows the ground truth evaluation of all viewpoints (note the UAV at the optimal position), and 4) the final result. Fig. 12 shows the result for a case with obstacles blocking the direct view of the facade of a target. With our method, the UAV captures two images from each side of the T-shaped obstacle to fully cover the front surface, and only requires one angled shot to inspect the entire side surface. Lastly, Fig. 13 shows the results for a setting with a slash-shaped obstacle. The ROSbot UGV locates the best vantage point at which the obstacle appears as a thin line.
@@
To further reinforce these results, we performed virtual experiments using a simulated RotorS Firefly equipped with a stereo camera. This experiment ran on the same hardware in Sec. V in C++. The evaluation of each candidate viewpoint occurs in 2.39ms. As Fig. 14 shows the results of an example virtual experiment, the vehicle detects the tree in front of the house using the stereo camera and stores the readings in an OctoMap [24]. The vehicle uses PSO to locate the optimal viewpoint. The picture in Fig. 14(b) shows the resulting image from this viewpoint. For this experiment, we set $\beta$ to 0.8 and use RRT to route the vehicle around obstacles. The Gazebo snapshot in Fig. 14(c) shows the planned trajectory of the UAV. The Rviz snapshot in Fig. 14(d) shows the resulting plan given the observed obstacles, where the red-shaded region represents the target.
Figure 14: An example virtual experiment where a tree obstructs the view of the target. (a) and (b) show the initial and final view. (c) and (d) show the trajectory followed to the best viewpoint in the gazebo and rviz.